OpenStudy (anonymous):

what value of k will x+1 be a factor of x^13-2x^7+3x+k? do i have to divide it out? or is there an easier way?

OpenStudy (anonymous):

synthetic division maybe

OpenStudy (anonymous):

or actually no there is an easier way

OpenStudy (dan815):

sub -1 and see wat k works

OpenStudy (anonymous):

if \(x+1\) is a factor, then \(-1\) is a zero replace \(x\) by \(-1\) set the result equal to zero, solve for \(k\)

OpenStudy (anonymous):

what @dan815 said

OpenStudy (anonymous):

so um x^13-2x^7+3x+k=0 and x is -1 right!?

OpenStudy (dan815):

hey i think u must understand what a factor really means

OpenStudy (dan815):

basically for x+1 to be a factor it means (x+1)*(something)= x^13-2x^7+3x+k

OpenStudy (dan815):

so if x=-1 then (x+1) this bracket will be = 0 so the whole expression will be 0 and similarily the right side = 0

OpenStudy (dan815):

so yes now you can say when -1 then x^13-2x^7+3x+k =0 and see which value for k works

OpenStudy (anonymous):

oh yeah alright! i understand now! thanks a lot!

OpenStudy (dan815):

yep you are welcome! see you